# Breaking the SFT Plateau: Multimodal Structured Reinforcement Learning for Chart-to-Code Generation

### v1

@@ 4.1 Baselines and Benchmarks
We evaluate MSRL against three categories of state-of-the-art models. The first category comprises general-domain, open-source Multimodal Large Language Models (MLLMs): InternVL2 (8B, 26B) Chen et al. (2024), Qwen2-VL (7B, 72B) Wang et al. (2024), and Qwen2.5-VL-7B Bai et al. (2025). The second includes proprietary models: GeminiProVision Team et al. (2023), Claude-3-opus Anthropic (2024), GPT-4V OpenAI (2023), and GPT-4o OpenAI (2024). The third category consists of chart-specific MLLMs: ChartLlama Han et al. (2023), Tinychart Zhang et al. (2024b), ChartVLM Xia et al. (2024), Chart2Code Zhang et al. (2025) and ChartCoder Zhao et al. (2025).
All models were evaluated in a zero-shot setting on two distinct benchmarks: ChartMimic Shi et al. (2024) and ReachQA He et al. (2024). For ReachQA, we adopt the evaluation setting of Chart2Code Zhang et al. (2025), using the 500 plotting scripts from its test set. For ChartMimic, we utilize the 600 examples from the latest version of the Direct Mimic task.
@@
Table 1: Evaluation results of various baseline models. $\ddagger$} denotes the updated result with code_pass constraints reported in its repository.
4.2 Implementation Details
We utilize 2.8M chart2code data for SFT and 33k curated high-quality chart2code data for RL. MSRL-SFT employs Qwen2.5VL-7B-Instruct as the baseline model and LLaMA-Factory Zheng et al. (2024) for SFT. The SFT stage is trained for one epoch with a learning rate of 1e-5 and a batch size of 32. The training process takes 60 hours on 16 H800 GPUs. Built upon the SFT model, MSRL undergoes a two-stage RL process using the MM-EUREKA framework Meng et al. (2025a). MSRL consists of two RL stages. Stage one utilizes 22k data pairs and a textual-only reward ($w_{t}=1$, $w_{v}=0$, $w_{e}=0.5$). Stage two uses the remaining 11k pairs with a hybrid reward ($w_{t}=0.5$, $w_{v}=0.5$, $w_{e}=0.5$). The textual reward components are weighted as $w_{d}=0.4$, $w_{c}=0.3$, $w_{l}=0.1$, $w_{t}=0.1$, $w_{lbl}=0.1$. We deploy Qwen2.5-VL-72B-Instruct as the evaluation model with max_model_len set to 12,288 for calculating visual rewards. For both RL stages, we train for one episode with a learning rate of 1e-6 and a batch size of 128. We set the generation temperature to 1, generate 8 rollouts per sample, and omit the KL divergence from the loss computation. The first and second stages required 10 and 24 hours of training, respectively, on 24 H800 GPUs.
Model ChartMimic ReachQA
@@ 4.2 Implementation Details
GPT-4o 81.5 89.8 77.3 67.2 84.4 91.2 81.3 70.5
Qwen2-VL-72B 43.2 80.5 54.6 39.4 41.0 52.2 59.1 47.8
Qwen2.5-VL-7B 37.7 65.6 42.6 32.4 29.6 40.9 44.2 32.8
ChartCoder 65.9 85.4 72.3 66.4 58.2 80.0 70.4 63.2
MSRL 78.0 93.0 75.3 68.2 87.5 96.6 83.1 77.2
@@ 4.3 Main Results
Our proposed MSRL establishes a new state-of-the-art among all open-source MLLMs, as detailed in Table 1. On the ChartMimic benchmark Shi et al. (2024), it achieves an execution rate of 96.5, a low-level score of 78.6, and a high-level score of 83.8. Notably, its performance is not only superior to its open-source models but also comparable to that of GPT-4o, a significantly larger proprietary model. On the ChartMimic benchmark, MSRL outperforms ChartCoder, the previous leading Chart-domain MLLM specifically designed for chart-to-code generation. Unlike other works that yield suboptimal performances from fine-tuning open-source models, our initial MSRL-SFT model already establishes a new state-of-the-art among open-source models. However, MSRL-SFT still falls short of GPT-4o on both low-level and high-level performance metrics. Building upon this strong foundation, the MSRL strategy further boosts the performance, especially in high-level scores, demonstrating the effectiveness of our integrated approach. The performance gap between SFT and RL models is emphasized in Figure 1.
@@
SFT Plateau To investigate and demonstrate the effectiveness and efficiency of RL, we utilize various data scales for SFT to explore the upper limit of it. As illustrated in Figure 1, the results reveal a clear trend, whereby model performance improves sharply as the data size increases to 400k samples, followed by a period of slower growth up to the 1.6M samples. Beyond this data scale, performance saturates and reaches a distinct plateau, showing negligible growth with any further increase in data scale. Based on this trend, even if the data is multiplied, the benefits brought by SFT will be minimal. Applying our proposed MSRL method yields significant performance gains across all metrics. On the ChartMimic benchmark, we observe an average improvement of 5%, with notable increases in execution rate as well as in both low-level and high-level scores.
@@ 4.4 Ablation Study
In this subsection, we perform various ablation studies to explore and verify the proposed two-stage training stages and multi-granularity reward functions.
@@
Figure 4: Comparison of reward gains during RL training with various reward settings.
Rewards ChartMimic
Textual Visual Exec.Rate Low-Level High-Level
93.17 73.02 77.61
97.00 78.06 82.69
96.83 77.80 81.90
96.33 78.56 82.44
Table 4: Ablation studies on various reward functions.
Reward Ablation To systematically evaluate the effectiveness of our proposed multi-granularity reward systems, we conduct ablation studies on various reward functions. To isolate the impact of the reward functions from our two-stage RL strategy, all experiments in Table 4 are trained using only a single stage. The results show that merely textual reward and combining textual and visual rewards both improve the performance in both low-level and high-level scores to a new SOTA. Also, the result demonstrates that directly utilizing visual rewards on the SFT model is not as effective as the textual reward. The detailed training process is illustrated in Figure 4. However, both of them show significant performance improvement compared with the SFT model, which leads us to explore the two-stage RL training strategy.
Stage Ablation. We also conduct an ablation study on our proposed two-stage RL training strategy. In this study, we compare three different reward settings: (1) a combination of textual and visual rewards in a single stage; (2) textual rewards in the first stage and visual rewards in the second; and (3) textual rewards in the first stage followed by both textual and visual rewards in the second. As shown in Table 5, using textual rewards in the first stage followed by combined textual and visual rewards in the second yields the best performance.
Figure 5: Showcasing charts generated by MSRL compared to proprietary and open-source MLLMs. The charts produced by MSRL align well with their ground-truth ones.
RL Stage ChartMimic
Stage 1 Stage 2 Exec.Rate Low-Level High-Level
Textual + Visual 96.33 78.56 82.44
Textual Visual 96.17 77.29 82.61
Textual Textual + Visual 96.50 78.62 83.83
Table 5: The ablation study of the reward functions on various RL stages.
@@ 5 Conclusion
In this work, we investigate the performance plateau of Supervised Fine-Tuning (SFT) in chart-to-code generation and propose a Multimodal Structured Reinforcement Learning (MSRL) strategy to overcome it. We begin by constructing a 3M-pair training corpus from real-world tables to systematically demonstrate that merely increasing SFT data leads to diminishing returns. To overcome this limitation, we introduce the MSRL strategy. It optimizes the saturated SFT model through a two-stage training process, guided by a novel multi-granularity reward function that assesses both textual and visual correctness. The textual reward employs a rule-based method to verify code accuracy across five key aspects, while the visual reward utilizes a render-and-compare mechanism, where an MLLM scores the similarity between the generated chart and the ground truth. Our experiments demonstrate that our model achieves new state-of-the-art performance on standard benchmarks, surpassing all open-source models and rivaling powerful proprietary ones. We make significant contributions to the field, including the curation of a large-scale dataset for a definitive empirical analysis of the SFT performance bottleneck, and the development of the MSRL framework.

### v2

@@ 4.1 Baselines and Benchmarks
We evaluate MSRL against three categories of state-of-the-art models. The first category comprises general-domain, open-source Multimodal Large Language Models (MLLMs): InternVL2 (8B, 26B) Chen et al. (2024b), Qwen2-VL (7B, 72B) Wang et al. (2024), and Qwen2.5-VL-7B Bai et al. (2025). The second includes proprietary models: GeminiProVision Team et al. (2023), Claude-3-opus Anthropic (2024), GPT-4V OpenAI (2023), and GPT-4o OpenAI (2024). The third category consists of chart-specific MLLMs: ChartLlama Han et al. (2023), Tinychart Zhang et al. (2024b), ChartVLM Xia et al. (2024), Chart2Code Zhang et al. (2025) and ChartCoder Zhao et al. (2025b).
All models were evaluated in a zero-shot setting on two distinct benchmarks: ChartMimic Yang et al. (2024a) and ReachQA He et al. (2024). For ReachQA, we adopt the evaluation setting of Chart2Code Zhang et al. (2025), using the 500 plotting scripts from its test set. For ChartMimic, we utilize the 600 examples from the latest version of the Direct Mimic task.
Table 2: Evaluation results of various models on the ChartMimic Direct Mimic and ReachQA benchmarks. $\ddagger$} denotes the updated result with code_pass constraints reported in its repository.
@@
Model Params ChartMimic ReachQA
@@
GPT-4o - 81.5 89.8 77.3 67.2 84.4 91.2 81.3 70.5
Qwen2-VL-72B 72B 43.2 80.5 54.6 39.4 41.0 52.2 59.1 47.8
Qwen2.5-VL-7B 7B 37.7 65.6 42.6 32.4 29.6 40.9 44.2 32.8
ChartCoder 7B 65.9 85.4 72.3 66.4 58.2 80.0 70.4 63.2
MSRL 7B 78.0 93.0 75.3 68.2 87.5 96.6 83.1 77.2
@@ 4.2 Main Results
Our proposed MSRL establishes a new state-of-the-art among all open-source MLLMs, as detailed in Table 2. On the ChartMimic benchmark Yang et al. (2024a), it achieves an execution rate of 96.5, a low-level score of 78.6, and a high-level score of 83.8. Notably, its performance is not only superior to its open-source models but also comparable to that of GPT-4o, a significantly larger proprietary model. On the ChartMimic benchmark, MSRL outperforms ChartCoder, the previous leading Chart-domain MLLM specifically designed for chart-to-code generation. Unlike other works that yield suboptimal performances from fine-tuning open-source models, our initial MSRL-SFT model already establishes a new state-of-the-art among open-source models. However, MSRL-SFT still falls short of GPT-4o on both low-level and high-level performance metrics. Building upon this strong foundation, the MSRL strategy further boosts the performance, especially in high-level scores, demonstrating the effectiveness of our integrated approach. The performance gap between SFT and RL models is emphasized in Figure 1.
@@
SFT Plateau To investigate and demonstrate the effectiveness and efficiency of RL, we utilize various data scales for SFT to explore the upper limit of it. As illustrated in Figure 1, the results reveal a clear trend, whereby model performance improves sharply as the data size increases to 400k samples, followed by a period of slower growth up to the 1.6M samples. Beyond this data scale, performance saturates and reaches a distinct plateau, showing negligible growth with any further increase in data scale. Based on this trend, even if the data is multiplied, the benefits brought by SFT will be minimal. Applying our proposed MSRL method yields significant performance gains across all metrics. On the ChartMimic benchmark, we observe an average improvement of 6%, with notable increases in execution rate as well as in both low-level and high-level scores.
Figure 4: Performance curve of MSRL with increasing data scale. From left to right, the curves correspond to the SFT stage, RL stage 1, and RL stage 2. Avg Score denotes the average of the Low- and High-level scores in ChartMimic.
RL Plateau To investigate whether RL training exhibits a performance plateau, we present the performance curves of the two-stage RL training with the increase of data size, as shown in Figure 4. In the first stage using only textual rewards, the model breaks through the SFT plateau and achieves a significant performance improvement of approximately 5%, converging at a maximum of 22k training samples. In the second stage with multimodal rewards, the model’s performance is further improved by about 1.5%, and almost converges at a maximum of 11k training samples. Our experiments effectively capture the performance plateau and improvements within a practical data size range, providing strong evidence for the effectiveness of our two-stage RL strategy.
@@ 4.3 Ablation Study
Table 4: Ablation studies on SFT and RL stages. For two RL experiments, only a textual reward is used.
@@
Table 5: Comparison of single-stage and two-stage RL training with different reward strategies. “T” and “V” represent textual and visual reward, respectively. All training is conducted on H800 GPUs.
Rewards Samples ChartMimic GPU Hours
Exec.Rate Low-Level High-Level
- - 93.17 73.02 77.61 -
(T, -) (22k, -) 97.00 78.06 82.69 240
(V, -) (22k, -) 97.37 79.13 84.01 1344
(T, T + V) (22k, 11k) 97.50 79.50 83.81 912
(T, T + V) (22k, 5.5k) 96.50 78.62 83.83 576
Reward and RL Training Strategy Ablation To thoroughly evaluate the effects of reward design and RL training strategies, we conduct ablation experiments comparing textual rewards, visual rewards, and their combinations. As shown in Table 5, both reward types improve model performance, but differ notably in resource consumption and efficiency. Visual rewards provide notable enhancements in code details and visual fidelity, but require more computational resources due to the overhead of image rendering and model evaluation. In contrast, textual rewards achieve competitive performance with much lower resource requirements. To balance efficiency and effectiveness, we adopt a two-stage RL strategy. The ablation results yield several important insights: 1) Both textual and visual rewards lead to significant improvements over the SFT model, with visual rewards being more effective for visual fidelity. 2) The high resource consumption of visual rewards highlights the importance of balancing performance and efficiency. 3) The two-stage RL strategy achieves the trade-off by first using low-cost textual rewards, followed by high-cost visual rewards with a limited number of samples. Based on these findings, we empirically report the two-stage training strategy with 50% of the training samples in our experiments, as it offers a better trade-off between performance and consumption efficiency.
Generalization to Unseen Plotting Libraries To evaluate whether MSRL can transfer its capabilities to other plotting libraries, we collect 107 Seaborn and 150 Plotly images from their official websites. Since we only use Matplotlib-style code to generate charts, the new test data are out-of-domain. We perform inference with both Qwen2.5-VL-7B and MSRL on these two small test sets, calculating execution rates and high-level scores according to the ChartMimic protocol. As shown in Table 6, the results demonstrate that MSRL exhibits generalization to unseen Seaborn- and Plotly-style images.
Table 6: Performance of Qwen2.5-VL-7B and MSRL on Seaborn and Plotly test sets.
Model Seaborn Plotly
Exec.Rate High-Level Exec.Rate High-Level
Qwen2.5-VL-7B 69.2 25.7 62.7 22.6
MSRL 85.1 30.5 90.0 35.9
Figure 5: Showcasing charts generated by MSRL compared to proprietary and open-source MLLMs. The first column shows the ground-truth chart images. Columns 2–4 display charts rendered from the code generated by GPT-4o, ChartCoder, and Qwen2.5-VL-7B, while the last column shows charts rendered by our MSRL model.
@@ 5 Conclusion
In this work, we investigate the performance plateau of Supervised Fine-Tuning (SFT) in chart-to-code generation and propose a Multimodal Structured Reinforcement Learning (MSRL) strategy to overcome it. We begin by constructing a 3M-pair training corpus from real-world tables to systematically demonstrate that merely increasing SFT data leads to diminishing returns. To address this, we introduce the MSRL, which optimizes the saturated SFT model using a two-stage training process guided by a multi-granularity reward function that assesses both textual and visual correctness. Our approach achieves new state-of-the-art performance on standard benchmarks, surpassing all open-source models and rivaling powerful proprietary ones. We make significant contributions to the field, including the curation of a large-scale dataset for a definitive empirical analysis of the SFT performance bottleneck, and the development of the MSRL framework.
6 Reproducibility Statement
For datasets, we provide a detailed description of the data generation process in Section 3.1 of the paper, along with dataset statistics and examples in Appendix A, and the exact prompts used for dataset construction in Appendix E. For code implementation, we include the SFT and RL training code for MSRL in the supplementary materials, which contains complete training scripts and implementations of the multimodal reward functions.
Appendix A Chart2Code Details
Table A: The detailed chart types and corresponding quantities used in the SFT and RL phases.
Split Bar Line ErrorBar Heatmap Box Scatter Histogram Radar 3D
SFT 131,623 131,596 130,818 123,795 125,541 127,612 132,180 124,160 128,744
RL 2,669 2,743 1,067 1,984 756 2,639 1,510 776 1,699
Split Pie ErrorPoint Violin Area Bubble Multi-axes Ring Rose Treemap
SFT 125,812 123,774 120,180 120,038 127,207 116,438 105,912 118,047 119,176
RL 1,446 2,277 900 1,558 1,093 1,691 960 674 518
Split Bar_num Contour Density Quiver Graph Funnel Total
SFT 128,333 123,461 130,769 128,987 108,150 1,152 2,853,505
RL 2,757 1,040 1,899 1,700 - - 34,356
We count the quantity and distribution of 24 chart types used in the SFT and RL stages respectively, as detailed in Table A. The distribution among the various types is relatively balanced. For the RL data, we filtered out two chart types, Graph and Funnel, which are not suitable for textual reward calculation. Furthermore, we present three visualization examples from the Chart2Code dataset, with increasing data density and complexity, as shown in Figures G, H and I.
Appendix B Implementation Details
We utilize 2.8M chart2code data for SFT and 33k curated high-quality chart2code data for RL. MSRL-SFT employs Qwen2.5VL-7B-Instruct as the baseline model and LLaMA-Factory Zheng et al. (2024) for SFT. The SFT stage is trained for one epoch with a learning rate of 1e-5 and a batch size of 32. The training process takes 60 hours on 16 H800 GPUs. Built upon the SFT model, MSRL undergoes a two-stage RL process using the MM-EUREKA framework Meng et al. (2025). MSRL consists of two RL stages. Stage one utilizes 22k data pairs and a textual-only reward ($w_{t}=1$, $w_{v}=0$, $w_{e}=0.5$). Stage two uses the remaining 11k pairs with a hybrid reward ($w_{t}=0.5$, $w_{v}=0.5$, $w_{e}=0.5$). The textual reward components are weighted as $w_{d}=0.4$, $w_{c}=0.3$, $w_{l}=0.1$, $w_{t}=0.1$, $w_{lbl}=0.1$. We deploy Qwen2.5-VL-72B-Instruct as the evaluation model with max_model_len set to 12,288 for calculating visual rewards. For both RL stages, we train for one episode with a learning rate of 1e-6 and a batch size of 128. We set the generation temperature to 1, generate 8 rollouts per sample, and omit the KL divergence from the loss computation. The first and second stages required 10 and 24 hours of training, respectively, on 24 H800 GPUs.
Appendix C More Ablation Studies
The Use of Visual Evaluator To verify that the improvement in performance results from effective visual reward design rather than overfitting to the preferences of the VLLM judger, we conduct an ablation experiment to observe the distribution of sample scores when the visual evaluator changes. Specifically, we randomly sample 200 instances from the RL data and use the MSRL-SFT model to perform chart-to-code prediction, rendering the resulting Python code into images. Following the ChartMimic scoring guidelines, human ratings of rendered samples serve as the gold scores. Next, we use GPT-4.1, Qwen2.5-VL-72B, Qwen2.5-VL-7B, and InternVL3-38B to score the images. Additionally, we compute the visual similarity between the ground-truth and rendered images using DINO-L. We calculate the Pearson correlation coefficients between the scores from these models and the human scores, and plot the distributions of model scores versus human scores. As shown in the Figure A, GPT-4.1 achieves the highest correlation with human judgment and has the most similar score distribution, followed by Qwen2.5-VL-72B, while Qwen2.5-VL-7B performs the worst. Considering both performance and budget, we select Qwen2.5-VL-72B as the visual evaluator, as it fairly reflects model capability and prevents reward hacking.
In addition, we introduce two vision models, InternViT Chen et al. (2024a) and DINOv2-L Oquab et al. (2023), as evaluators for visual rewards, calculating the visual similarity between ground-truth and generated image features. As shown in Table B, these vision models yield improved performance, but there remains a substantial gap compared to models employing VLMs as visual evaluators. This indicates that vision models trained on image-text matching assess similarity at a coarse level and do not adequately capture fine-grained differences in chart elements. Compared to Qwen2.5-VL-72B, smaller vision models such as DINO-L exhibit greater discrepancies in score distributions relative to human judgment.
Figure A: Scatter plots of model scores versus human gold scores for five visual evaluators (GPT-4.1, Qwen2.5-VL-72B, Qwen2.5-VL-7B, InternVL3-38B, and DINO-L). Pearson $r$ denotes the Pearson correlation coefficient between model and human scores.
Table B: The ablation study of visual evaluators. ”-” denotes the SFT model. All experiments exclude textual rewards.
Visual Evaluator ChartMimic
Low-Level High-Level
- 73.0 77.6
InternViT 75.1 78.1
DINOv2-L 75.5 79.8
Qwen2.5-VL-72B 77.8 81.9
GRPO vs DPO Intuitively, GRPO is naturally more suitable for chart-to-code tasks than DPO. The degree to which code restores a chart can be quantitatively measured by a set of well-defined rules, making it ideal for designing the reward function in GRPO. In contrast, DPO learns from preference pairs, and it is challenging to clearly define what constitutes ”bad” code in chart-to-code tasks, which can lead to ambiguity in the model’s optimization direction. To ensure a thorough evaluation, we conduct an ablation study comparing GRPO and DPO. Since DPO requires the construction of positive and negative sample pairs, we follow the Chart2Code Zhang et al. (2025) and randomly perturb 2-3 attributes of the chart code to generate negative samples using GPT-4.1, resulting in an RL dataset of comparable size to GRPO for DPO training on the SFT model. As shown in Table C, the performance improvement of DPO is significantly less than that of GRPO. This indicates both the superiority of the GRPO algorithm and the effectiveness of our multimodal structured reward design and two-stage curriculum learning.
Table C: The ablation study of RL optimizers. ”-” denotes the SFT model.
RL Optimizer ChartMimic
Low-Level High-Level
- 73.0 77.6
DPO 74.0 78.9
GRPO 78.6 83.8
Comparison with Existing Datasets To highlight the value of our synthetic data, we compare the results of using our data and open-source data (i.e., ChartCoder) in both the SFT and RL stages, as shown in Table D. The results show that models trained with our data achieve better performance in both stages. On the one hand, our sufficient data enable SFT to reach its full potential. On the other hand, our charts are generated from real tabular data, featuring more challenging multi-chart content and greater visual diversity, which are key factors contributing to the effectiveness of our dataset.
Table D: Performance comparison with the ChartCoder dataset. We use Qwen2.5-VL-7B and MSRL-SFT as base models for SFT and RL for fine-tuning on the two datasets, respectively.
Stage Dataset ChartMimic
Low-Level High-Level
SFT ChartCoder 68.5 68.0
Ours 73.0 77.6
RL ChartCoder 76.5 81.2
Ours 78.6 83.8
Appendix D Qualitative Analysis
Figures E and  F show two examples of generated code from different models and their corresponding rendered images. Compared to the initial Qwen2.5-VL-7B model, both MSRL-SFT and MSRL show significant improvements in chart type recognition, numerical extraction, and color identification. Compared to MSRL-SFT, MSRL performs better in chart type recognition and clarity. In Figure E, MSRL-SFT incorrectly identifies the line chart as a scatter plot, displays incorrect legends, and has overlapping issues with image boundaries. In Figure F, MSRL-SFT has layout syntax errors, where only the last subplot is successfully displayed out of three subplots. MSRL’s results are more consistent with the original images, though it still has some minor issues in detail. For example, in Figure E, MSRL arbitrarily adds a title and scatter data points to the subcategory “Ours (LoSA)”. In Figure F, while MSRL successfully renders three subplots, the layout recognition is not sufficiently accurate, leaving room for further improvement.
Appendix E Prompt Template
To enhance transparency and reproducibility, we provide the exact prompts used for dataset generation and visual reward feedback.
Figure B shows the prompt used for code generation. We use real table data as input, select one chart type from 24 predefined chart types, and sample code examples of the selected chart type to generate plotting code. Specifically, for chart types Pie, Ring, and Treemap, we require the model to display actual values on the image, as these chart types typically only visualize percentage values in their code, making it impossible to read the true values from the image.
Figure C shows the prompt used for evaluating image quality. When processing RL data, we use this prompt to filter out visually low-quality samples, such as those with overlapping text or elements.
Figure D shows the prompt used to calculate visual rewards. Following ChartMimic Yang et al. (2024a), we assess the match between ground truth and model predictions across 6 dimensions, with this score normalized as the visual score and incorporated into the overall reward computation.
Figure B: Prompt template for Chart2Code code generation.
Figure C: Prompt template for image-quality evaluation during RL data filtering.
Figure D: Prompt template for visual reward feedback during the RL training.
Appendix F The Use of Large Language Models (LLMs)
In this work, we utilized LLMs primarily as writing assistance tools to polish language and improve clarity. They helped refine sentence structures and enhance readability throughout the manuscript. The core research ideas, experimental design, implementation, and analysis were all conducted by the authors, with LLMs serving only as editorial aids.
Figure E: An example of comparing the code generated by different models with rendered images.
Figure F: An example of comparing the code generated by different models with rendered images.
Figure G: A case of chart-code pairs for area charts in our Chart2Code dataset.
Figure H: A case of chart-code pairs for errorbar charts in our Chart2Code dataset.
Figure I: A case of chart-code pairs for bar_num charts in our Chart2Code dataset.